OpenStudy (anonymous):

A certain office building infrequently experiences power outages. It is known that the probability of a power outage on any given day is 0.005, and power outages are independent of each other. a) What is the probability that during any given 400 day period there will be one day with a power outage? b) What is the probability that power outages will occur on more than three days during the 400 day period?

OpenStudy (anonymous):

a) 400*.005 b) 400*.005*.005*.005 (Not sure on this one)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@whpalmer4

OpenStudy (perl):

a) 400 *.005

OpenStudy (perl):

b) Let's define a variable X = # of days of power outtage in 400 days. so X can equal 0,1,2,3... 400 we want probability P ( X > 3 )

OpenStudy (perl):

Now the complement is much easier since P(X>3) = P( X = 4) + P(X=5) + ... P(X = 400). But P(X>3) = 1- P(X<=3 ) = 1 - ( P(x=0) + P(x=1) +P(x=2) +P(x=3) )

OpenStudy (anonymous):

I did part a like this == pr= (1/400)*0.005

OpenStudy (perl):

it should be, n*p

OpenStudy (perl):

400 * .005

OpenStudy (anonymous):

@perl 1-(1/400+2/399+3/398)x0.005 for part B

OpenStudy (perl):

this is a binomial probability problem P(S) = .995 P(F) = .005 we want 1 - [ P( X = 0 ) + P( X = 1) + P( X = 2) + P( X = 3) ] 1 - ( .995^400 + (400 choose 1) * .995^399 * .005 + (400 Choose 2)*.995^398*.005^2+ (400 choose 3) * .995^397 * .005^3 )

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What are you tagging me for when @perl is right there helping you?

OpenStudy (anonymous):

it dose not make sense , cuz he is multiply success and failure prpbabilities with each other in same time ,

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

I have to leave, perhaps @mathmale can help you.

OpenStudy (perl):

kalyan, this is a similar question http://users.rowan.edu/~schultzl/TI/binomial_StatI.pdf

OpenStudy (perl):

sorry i dont understand your question

OpenStudy (perl):

b) 1 - binomcdf( 400, .005, 3) = .1424